concentration of Ag+ ions in saturated solution of Ag2CrO4  at 200C is 1.5 * 10-4 mol L-1 . At 200C, the solubility product of Ag2CrO4 is-

Grade:

1 Answers

Arun
5 years ago
Dear Tanusri
 
Ksp = [Ag+] = 1. 5 * 10^-4
 
therefore [C2O4]2- = 1.5 * 10^-4/ 2 = 0.75 * 10^-4
 
Ksp = (1. 5 *10^-4 )^2 * (0.75 * 10^-4))
 
 = 1.6875 * 10^-12